OpenStudy (anonymous):

A barrel contained 10 gallons of water and 2 gallons of oil. How many gallons of a solution that is 5% oil and 95% water must be added to the barrel to dilute the oil down to 10% of the total volume in the barrel?

OpenStudy (paxpolaris):

oil and water don't mix

OpenStudy (anonymous):

lol. I'm terrible at these kinds of problems for some reason...can you guide me through it?

OpenStudy (paxpolaris):

if you are adding x gallons of solution:\[\Large {2+0.05x \over 12+x} = 10\%\]

OpenStudy (anonymous):

? I don't really get what that means

OpenStudy (paxpolaris):

after adding x gallons of the 5% oil solution: total volume = x plus the 12 gallons already in barrel volume of oil= 2gallons alreasy in barrel + 5% of x added

OpenStudy (paxpolaris):

after adding solution \[\Large {\text{volume of oil} \over \text{total volume}} =10\% \ \ or\ \frac1{10}\]

OpenStudy (anonymous):

x + 12 = y 0.05x + 2 = 0.1y 0.05x + 2 = 0.1(x + 12) 0.05x + 2 = 0.1x + 1.2 0.05x = 0.8 x = 16 Did I do it correctly?

OpenStudy (paxpolaris):

never mind it 's correct

OpenStudy (anonymous):

Alright. Thanks! I can never seem to do these problems. Is there a different way that you would do it or is that the way you're supposed to do it?

OpenStudy (paxpolaris):

all ways are pretty similar....

OpenStudy (anonymous):

Oh alright THanks :)